# Interface-sensitive microwave loss in superconducting tantalum films sputtered on c-plane sapphire

### v1

@@ Abstract
Quantum coherence in superconducting circuits has increased steadily over the last decades as a result of a growing understanding of the various loss mechanisms. Recently, tantalum (Ta) emerged as a promising material to address microscopic sources of loss found on niobium (Nb) or aluminum (Al) surfaces. However, the effects of film and interface microstructure on low-temperature microwave loss are still not well understood. Here we present a systematic study of the structural and electrical properties of Ta and Nb films sputtered on c-plane sapphire at varying growth temperatures. As growth temperature is increased, our results show that the onset of epitaxial growth of -phase Ta correlates with lower Ta surface roughness, higher critical temperature, and higher residual resistivity ratio, but surprisingly also correlates with a significant increase in loss at microwave frequency. We determine that the source of loss is located at the Ta/sapphire interface and show that it can be fully mitigated by either growing a thin, epitaxial Nb inter-layer between the Ta film and the substrate or by intentionally treating the sapphire surface with in-situ argon plasma before Ta growth. In addition to elucidating this interfacial microwave loss, this work provides adequate process details that should allow for the reproducible growth of low-loss Ta film across fabrication facilities.
@@ I Introduction
In the field of superconducting quantum computing, the transmon -a Josephson junction (JJ) shunted by a capacitorKoch et al. (2007); Nakamura, Pashkin, and Tsai (1999)- is a promising candidate moving towards useful quantum computationKrantz et al. (2019); Kjaergaard et al. (2020). This technology has advanced rapidly, as qubit coherence times have benefited from improvements in measurement setup, device design layoutKrantz et al. (2019); Kjaergaard et al. (2020), and more recently advances in materials used to form the capacitor metalPlace et al. (2021); Bal et al. (2024). The capacitor surface and its interface with the substrate are known to host two-level systems (TLS) which act as dissipation channels which limits coherence times in superconducting qubitsSimmonds et al. (2004); Siddiqi (2021).
Tantalum (Ta) grown on widely available c-plane sapphire substrates has emerged as an especially promising candidate for low-loss superconductor-based capacitors. Ta was shown to offer coherence time improvements when compared to niobium (Nb) in reports of state-of-the-art transmons made using Ta-metal capacitor pads shunted by Al/AlOx/Al JJ’sPlace et al. (2021); Wang et al. (2022). The reason for the enhanced performance in Ta devices is thought to be due to the more favorable properties of the surface oxides of Ta as compared to Nb. Optimizing the growth and processing conditions of the capacitor material for low microwave (MW) loss and understanding the mechanisms of surface and interface loss is presently an active area of research.
Superconducting devices made from Ta typically utilize -phase Ta formed in the body-centered cubic, (BCC) structure, which has a higher superconducting critical temperature () of 4.49KInaba (1980) as compared to the tetragonal -phase Ta, which has a below 1KMazin (2022). For Ta thin films, room temperature deposition usually results in the metastable -phase, while formation of -phase Ta typically requires heating the substrate during growthJiang et al. (2005) or deposition on a seed layerLee and Windover (1998); Face and Prober (1987); Urade et al. (2024). This formation of two distinct elemental phases contrasts with Nb, which typically only forms thin films in the BCC structure. Notably, the lattice parameters of Ta and Nb in the BCC structure are nearly the same, having values of 0.3306 nm and 0.3301 nm respectivelyWard, Grier, and Petford-Long (2003).
@@
Increasing the size of the capacitor also increases the mode volume in the packaging environment, where care must be taken to avoid package related loss from lossy dielectrics and resistive (non-superconducting) metals in the package. In addition, resonant modes present in the MW environment of the package can also couple to the qubit and act as decoherence sources, limiting coherence times to below state-of-the-artHuang et al. (2021); Sheldon et al. (2017). Measurement of high coherence transmons also requires substantial shielding and filtering in the measurement setup to suppress the quasiparticle-generating radiation seen by the qubit, which may also limit coherence timesGordon et al. (2022).
In contrast, coplanar waveguide (CPW) resonators may also be used to assess MW loss arising from the superconducting metal layer and its interfaces while avoiding extra processing steps and measurement difficulties associated with transmon characterization. Resonator internal quality factor (), a figure of merit expressing the dissipation rate due to materials-related losses in the resonator, may be measured using a continuous wave measurement by fitting the reflected () or transmitted () resonance spectra to established modelsMcRae et al. (2020). Reducing the CPW resonator gap width confines the mode volume and increases the electric field participation at the metal surfaces and interfaces. Because MW loss is typically dominated by surface and interface contributions, reducing CPW gap width tends to reduce , while also reducing package-related contributions to MW loss and providing an exacting evaluation of the superconducting metal, surface, and interface.
In this work, the structural and electrical properties of Nb and Ta films grown on c-plane sapphire by sputtering have been studied as a function of growth temperature while adequate details are provided in order to reproduce these results in different tools and labs. The MW properties of the films were evaluated using narrow-gap CPW resonators, transport properties were measured using temperature dependent resistivity measurements, and the structure of the films was characterized using X-ray diffraction (XRD), atomic force microscopy (AFM), reflection high energy electron diffraction (RHEED), and transmission electron microscopy (TEM). Measurements of these films revealed an unknown source of MW loss in epitaxial Ta(111) films grown directly on sapphire. An experiment was devised to test the Ta/sapphire interface which suggests that the epitaxial -Ta(111)/Al2O3(0001) interface is a significant source of MW loss and that loss may be avoided by simple treatments to the substrate surface before growth.
@@ II.1 Materials Growth and Micro-fabrication
Wafers were secured to a silicon backing wafer and inconel sample holder using clips made from Ta wire. The silicon backing wafer was heated radiatively by a standard bank of quartz halogen IR lamps through a window in the inconel holder which was found to result in reproducible and uniform heating across the wafer surface. Variation between substrate surface temperature and measured thermocouple temperature can be significant, and is known to vary depending on details including the substrate holder and sample mounting techniqueMizutani (1988); Lee et al. (1991). For the sake of reproducibility across different labs, tools, and sample mounting configurations, the thermocouple temperature was calibrated to the sample surface temperature at the melting point of aluminum (T=660). This was done by depositing 500nm of aluminum on a bare sapphire wafer at room temperature and observing the phase transition by RHEED and optically through a viewport as the sample was slowly heated. The melting point of Al on the sapphire surface was found to occur at a thermocouple temperature of 790 establishing that the surface temperature is approximately 130 lower than the thermocouple reading around the Al melting point. Through the remainder of this work, "growth temperature" refers to the temperature read on the thermocouple, but it is important to note that the sample surface temperature is significantly lower.
@@
Following Ta or Nb deposition, the metal films were patterned into coplanar waveguide (CPW) resonators and Hall bars using standard photolithography and dry etching. The Ta or Nb coated wafer was first solvent cleaned to obtain a consistent starting surface, followed by spin coating of an HMDS-based primer and standard photoresist. The wafers were patterned using direct laser writing, descummed in an oxygen plasma, and etched in an ICP tool using the same Cl2/BCl3 etch chemistry for both Nb and TaBal et al. (2024). After etching, the photoresist was removed using an NMP-based remover held at 80 with ultrasonication followed by an isopropanol rinse. The finished wafer was again coated in photoresist to protect the surface during dicing. 7.5mm x 7.5mm diced chips were cleaned in NMP-based remover at 80 with ultrasonication followed by an isopropanol rinse. CPW resonator chips were additionally cleaned in 6:1 buffered HF (BHF) for 2 minutes followed by a DI rinse immediately before packaging and wiring.
Notably, the substrate preparation procedure used for this study is minimal, avoiding the acid or plasma etching steps and furnace annealing commonly used with sapphire substratesDwikusuma, Saulys, and Kuech (2002); Vlasov et al. (2016); Place et al. (2021). Initially, samples were prepared by acid cleaning for 10 minutes in a stabilized formulation of sulfuric acid and hydrogen peroxide compounds held at 80 (Piranha etch). This step was not found to have an appreciable effect on the resulting film structure or electrical properties, presumably because the as-received wafers were already clean. However, chips that were prepared from the same as-received wafers by covering with protective photoresist before dicing were found to have surface contamination after the resist was stripped which affected film growth. For these chips, a Piranha-type etch was found to result in a clean starting growth surface. Piranha etching is a recommended step in order to reproduce the results presented here.
@@ II.3 Electrical Measurements
Characterization of Hall bars was performed in a helium-4 flow cryostat using a standard lock-in technique to obtain film sheet resistance as a function of temperature. Excitation currents of 1-100 A were used in a 4-point configuration. Residual resistivity ratio (RRR), a commonly used figure of merit for metallic thin films, is defined as the resistance at 300K divided by the resistance just above the superconducting transition temperature ().
Microwave loss of deposited films was studied using CPW resonators. Each CPW resonator chip consists of 8 inductively-coupled, quarter-wave resonators in hanger configuration off a central feedline similar to the design presented by Kopas et al.Kopas et al. (2022). The CPW dimensions of conductor/gap are 6m/3m and resonant frequencies are designed to fall between 5.5-6.5 GHz. The narrow conductor and gap widths are chosen to increase the electric field participation at surfaces and interfaces to study their contribution to microwave lossGao (2008); McRae et al. (2020).
Clean CPW resonator chips were wire-bonded with aluminum wire and packaged in a 2-port package constructed of Au-plated OFHC Cu for microwave measurements in a dilution refrigerator with a base temperature of 35mK. The attenuation and amplification setup used is similar to that presented by McRae et al.McRae et al. (2020). Microwave loss of the resonators was characterized using a vector network analyzer (VNA) by fitting the power dependent spectra of the resonators to the diameter correction method (DCM)Khalil et al. (2012); McRae et al. (2020) model to extract internal () and external () resonator quality factors. A Josephson parametric amplifierAumentado (2020) installed at the 35mK stage was used for all low-power measurements (below 1000 photon occupation in the cavities).
@@ III.1 Overview of Experiment
Temperature dependent resistivity measurements of the Ta films grown at high temperature suggest high quality films. of epitaxial Ta films were measured to be near reported bulk values with RRR > 30 for 100nm thick films.
@@
The same microfabrication process is used for all samples presented here and is shown to yield high quality CPW resonators for all Nb films, but only for Ta films grown at lower temperature. Ta films grown at higher temperature are epitaxial and high-quality structurally, but yield lossy resonators, seemingly at odds with structural and DC electrical characterization results.
In light of these observations, subsequent samples were fabricated to test the Ta/sapphire interface. The details of this experiment are provided in a subsequent section, and the results suggest that the epitaxial Ta/sapphire interface is the dominant source of microwave loss at both high and low power in the samples having low . While our experiments identify the interface as the source of loss and provide techniques shown to eliminate it, the microscopic mechanism of the loss is left as an open question.
@@ III.2 Temperature dependent growth of Nb and Ta on
Figure 1 shows XRD measurement results taken from Nb and Ta films grown at varying substrate temperature. Each peak indicates a crystalline phase oriented perpendicular to the substrate surface, along the growth direction. For Nb, growth at room temperature results in a mostly textured (having one dominant crystallographic orientation along the growth direction), (110) oriented film with a detectable amount of Nb(111). Non-specular XRD measurements (not included) show that the Nb(110) component is randomly oriented in-plane. For Nb films grown at elevated temperature, only epitaxial Nb(111) is observed with no notable changes to structure as the growth temperature is increased in the temperature range studied. Non-specular measurements on these films confirm the expected epitaxial relationshipMayer et al. (1992) with .
XRD results from the Ta films (Fig. 1(b)) are somewhat more complex owing to the tendency of Ta thin films to crystallize in both the BCC (-Ta) and tetragonal (-Ta) structures. Ta grown at room temperature results in a film consisting mostly of highly textured -Ta(001) with a detectable amount of -Ta. In addition to out-of-plane texture, the -Ta(001) was found to be textured in-plane. Non-specular measurements (not included) show three rotational domains of 4-fold symmetric -Ta(001) epitaxially aligned to the substrate, consistent with the 12-fold symmetric RHEED pattern (not included) observed immediately following growth. Ta deposition at 380 resulted in a biphasic film having both textured -Ta(001) and epitaxial -Ta(111) components, while growth at temperatures above 500 resulted in only epitaxial -Ta(111) having the same structural relationship to the substrate as the epitaxial Nb(111) described previously.
Sheet resistance at 300K, RRR, and were extracted from temperature dependent resistance measurements for all films and is included in Table 1. For Nb films, RRR increases with growth temperature as the film structure changes from textured Nb(110) to epi-Nb(111), but is somewhat reduced for the highest growth temperatures.
@@
For the textured -Ta(001) film grown at 20, resistivity increased with decreasing measurement temperature (RRR<1) and was below the base temperature of the He-4 cryostat used in the measurement. Temperature dependent transport of the epitaxial -Ta(111) samples was consistent with high quality films. A RRR value of 29 was measured for the -Ta(111) grown at 520 which increased to 47 for the film grown at 630. The of the -Ta(111) films range from 4.24K-4.35K, near reported values of pure, bulk Ta (4.49K)Inaba (1980).
AFM measurements were also performed on all of the films studied. While changes in morphology were observed with changing growth temperature, all films were found to have root-mean-square (RMS) roughness values below 2nm and no strong correlation to transport properties was observed. RMS roughness values for both Nb and Ta are included in Tables 1 and 2 respectively. For higher temperature growth, triangular features are observed in the surface morphology. For Nb, the triangular morphology was observed for both the samples grown at 520 and 630. Similar morphology was observed for Ta films grown above 700 which are not included in this study.
@@
Figure 2 shows of CPW resonators plotted against growth temperature for both Ta and Nb films. To obtain the data shown, spectra were measured for at least 3 resonators on each chip at varying input power corresponding to to below photons in the CPW cavities. values taken below 0.5 photon occupation (at least 3 per resonator) were then averaged together and plotted in Fig. 2. Thus, each data point shown includes multiple measurements of low-power from 3-8 different resonators measured from the same wafer, with the minimum, maximum, mean, and standard deviation indicated. These low-power data may be considered as figures of merit that are representative of the material and reflect the degree of MW loss at the ultra-low electric field intensity where transmons operate.
@@
Measurement results from the Ta resonators are more interesting and somewhat unexpected. Firstly, it was found that resonators from the -Ta films grown at room temperature are high performing, with values exceeding those made from Nb resonators grown using the same conditions. Resonators containing -Ta were also shifted to lower frequency, presumably due to kinetic inductance.

### v2

@@ Abstract
Quantum coherence in superconducting circuits has increased steadily over the last decades as a result of a growing understanding of the various loss mechanisms. Recently, tantalum (Ta) emerged as a promising material to address microscopic sources of loss found on niobium (Nb) or aluminum (Al) surfaces. However, the effects of film and interface microstructure on low-temperature microwave loss are still not well understood. Here we present a systematic study of the structural and electrical properties of Ta and Nb films sputtered on c-plane sapphire at varying growth temperatures. As growth temperature is increased, our results show that the onset of epitaxial growth of -phase Ta correlates with lower Ta surface roughness, higher critical temperature, and higher residual resistivity ratio, but surprisingly also correlates with a significant increase in loss at microwave frequency. Notably, this high level of loss is not observed in Nb films prepared in the same way and having very similar structure. By experimentally controlling the surface on which the Ta film is nucleated, we determine that the source of loss is only present in samples having an epitaxial Ta/sapphire interface and show that it is apparently mitigated by either growing a thin, epitaxial Nb inter-layer between the Ta film and the substrate or by intentionally treating, and effectively damaging, the sapphire surface with an in-situ argon plasma before Ta growth. In addition to elucidating this interfacial microwave loss, this work provides adequate process details to aid reproducible growth of low-loss Ta films across fabrication facilities.
@@ I Introduction
In the field of superconducting quantum computing, the transmon -a Josephson junction (JJ) shunted by a capacitorKoch et al. (2007); Nakamura, Pashkin, and Tsai (1999)- is a promising candidate moving towards useful quantum computationKrantz et al. (2019); Kjaergaard et al. (2020). This technology has advanced rapidly, as qubit coherence times have benefited from improvements in measurement setup, device design layoutKrantz et al. (2019); Kjaergaard et al. (2020), and more recently advances in materials used to form the capacitor metal, where state-of-the-art transmon coherence times have been demonstrated using aluminum (Al)Biznarova et al. (2024), niobium (Nb)Bal et al. (2024); Kono et al. (2024), and tantalum (Ta)Place et al. (2021). The capacitor surface and its interface with the substrate are known to host two-level systems (TLS) which act as dissipation channels which limit coherence times in superconducting qubitsSimmonds et al. (2004); Siddiqi (2021).
Tantalum grown on widely available c-plane sapphire substrates has emerged as an especially promising candidate for low-loss superconductor-based capacitors. Ta was shown to offer coherence time improvements when compared to Nb in reports of state-of-the-art transmons made using Ta-metal capacitor pads shunted by Al/AlOx/Al JJ’sPlace et al. (2021); Wang et al. (2022). The reason for the enhanced performance in Ta devices has been proposed to be due to the more favorable properties of the surface oxides of Ta as compared to NbPlace et al. (2021); Crowley et al. (2023). Optimizing the growth and processing conditions of the capacitor material for low microwave (MW) loss and understanding the mechanisms of surface and interface loss is presently an active area of research.
Superconducting devices made from Ta typically utilize -phase Ta formed in the body-centered cubic, (BCC) structure, which has a higher superconducting critical temperature () of 4.49KInaba (1980) as compared to the tetragonal -phase Ta, which has a below 1KMazin (2022). For Ta thin films, room temperature deposition usually results in the metastable -phase, while formation of -phase Ta typically requires heating the substrate during growthJiang et al. (2005) or deposition on a seed layerLee and Windover (1998); Face and Prober (1987); Urade et al. (2024); Alegria et al. (2023). This formation of two distinct elemental phases contrasts with Nb, for which metastable phases do exist, but thin films are most commonly found in the BCC structureLee et al. (2023); Kunzmann et al. (2024). Notably, the lattice parameters of Ta and Nb in the BCC structure are nearly the same, having values of 0.3306 nm and 0.3301 nm respectivelyWard, Grier, and Petford-Long (2003).
@@
Increasing the size of the capacitor also increases the mode volume in the packaging environment, where care must be taken to avoid package related loss from lossy dielectrics and resistive (non-superconducting) metals in the package. In addition, resonant modes present in the MW environment of the package can also couple to the qubit and act as decoherence sources, limiting coherence times to below state-of-the-artHuang et al. (2021); Sheldon et al. (2017). Measurement of high coherence transmons also requires substantial shielding and filtering in the measurement setup to suppress the quasiparticle-generating radiation seen by the qubit that also limits coherence timesGordon et al. (2022).
In contrast, coplanar waveguide (CPW) resonators may also be used to assess MW loss arising from the superconducting metal layer and its interfaces while avoiding extra processing steps, measurement difficulties associated with transmon characterization, and additional loss associated with Josephson elementsMcRae et al. (2020). Resonator internal quality factor (), a figure of merit expressing the dissipation rate due to materials and environmental-related losses in the resonator, may be measured using a continuous wave measurement by fitting the reflected () or transmitted () resonance spectra to established modelsMcRae et al. (2020). Reducing the CPW resonator gap width confines the mode volume and increases the electric field participation at the metal surfaces and interfaces. Because MW loss is typically dominated by surface and interface contributions, reducing CPW gap width tends to reduce , while also reducing package-related contributions to MW loss and providing an exacting evaluation of the superconducting metal, surface, and interface.
In this work, the structural and electrical properties of Nb and Ta films grown on c-plane sapphire by sputtering have been studied as a function of growth temperature while including sufficient details to provide a starting point for reproducibility. The MW properties of the films were evaluated using narrow-gap CPW resonators, transport properties were measured using temperature dependent resistivity measurements, and the structure of the films was characterized using X-ray diffraction (XRD), atomic force microscopy (AFM), reflection high energy electron diffraction (RHEED), and transmission electron microscopy (TEM).
Measurements of these films revealed an unknown source of MW loss in epitaxial Ta(111) films grown directly on sapphire. An experiment was devised to test the Ta/sapphire interface by forming inter-layers of either epitaxial Nb(111) or amorphized sapphire between the crystalline substrate and the Ta films, which suggests that the epitaxial - Ta(111)/ Al2O3(0001) interface is a significant source of MW loss and that loss may be avoided by simple treatments to the substrate surface before growth. While the suggested inter-layers have been employed in high performing superconductive circuits Urade et al. (2024); Alegria et al. (2023), a systematic evaluation of their effects on structural and electrical properties and a clear comparison to epitaxial -Ta(111)/Al2O3(0001) films has not yet been shown in the literature.
@@ II.1 Materials Growth and Micro-fabrication
Wafers were secured to a silicon backing wafer and inconel sample holder using clips made from Ta wire. The silicon backing wafer was heated radiatively by a standard bank of quartz halogen lamps through a window in the inconel holder which was found to result in reproducible and uniform heating across the wafer surface. Variation between substrate surface temperature and measured thermocouple temperature can be significant, and is known to vary depending on details including the substrate holder and sample mounting techniqueMizutani (1988); Lee et al. (1991). For the sake of reproducibility across different labs, tools, and sample mounting configurations, the thermocouple temperature was calibrated to the sample surface temperature at the melting point of aluminum (T=660). This was done by depositing 500nm of aluminum on a bare sapphire wafer at room temperature and observing the phase transition by RHEED and visually through a viewport as the sample was slowly heated. The melting point of Al on the sapphire surface was found to occur at a thermocouple temperature of 790 establishing that the surface temperature is approximately 130 lower than the thermocouple reading around the Al melting point. Through the remainder of this work, "growth temperature" refers to the temperature read on the thermocouple, but it is important to note that the sample surface temperature is significantly lower.
@@
Following Ta or Nb deposition, the metal films were patterned into coplanar waveguide (CPW) resonators and Hall bars using standard photolithography and dry etching. The Ta or Nb coated wafer was first solvent cleaned to obtain a consistent starting surface, followed by spin coating of a heximethyldisilazane (HMDS)-based primer and standard photoresist. The wafers were patterned using direct laser writing, descummed in an oxygen plasma, and etched in an inductively coupled plasma (ICP) tool using the same chlorine/boron trichloride (Cl2/BCl3) etch chemistry for both Nb and TaBal et al. (2024). After etching, the photoresist was removed using an n-methylpyrrolidone (NMP)-based remover held at 80 with ultrasonication followed by an isopropanol rinse. The finished wafer was again coated in photoresist to protect the surface during dicing. 7.5mm x 7.5mm diced chips were cleaned in NMP-based remover at 80 with ultrasonication followed by an isopropanol rinse. CPW resonator chips were additionally cleaned in 6:1 buffered hydrofluoric acid (BHF) for 2 minutes followed by a deionized water rinse immediately before packaging and wiring.
Notably, the substrate preparation procedure used for this study is minimal, avoiding the acid or plasma etching steps and furnace annealing commonly used with sapphire substrates Dwikusuma, Saulys, and Kuech (2002); Vlasov et al. (2016); Place et al. (2021). Initially, samples were prepared by acid cleaning for 10 minutes in a stabilized formulation of sulfuric acid and hydrogen peroxide compounds held at 80 (Piranha etch). This step was not found to have an appreciable effect on the resulting film structure or electrical properties, presumably because the as-received wafers were already clean. However, chips that were prepared from the same as-received wafers by covering with protective photoresist before dicing were found to have surface contamination after the resist was stripped which affected film growth. For these chips, a Piranha-type etch was found to result in a clean starting growth surface. Piranha etching is a recommended step in order to reproduce the results presented here. The samples studied in the main text of this work did not undergo a Piranha clean, though some additional results using Piranha cleaning are included in Appendix A.
@@ II.3 Electrical Measurements
Characterization of Hall bars was performed in a helium-4 flow cryostat. A standard lock-in configuration was used to obtain film sheet resistance as a function of temperature using the voltage output of the lock-in amplifier operated at 87 Hz in series with a bias resistor. The series combination of the device under test and the cryostat wiring is under , allowing the approximation of the voltage source/bias resistor combination as a current source. Excitation currents of 1-100 A were used in a 4-point configuration. Residual resistivity ratio (RRR), a commonly used figure of merit for metallic thin films, is defined as the resistance at 300K divided by the resistance just above the superconducting transition temperature ().
Microwave loss of deposited films was studied using CPW resonators. Each CPW resonator chip consists of 8 inductively coupled, quarter-wave resonators in hanger configuration off a central feedline similar to the design presented by Kopas et al. Kopas et al. (2022). The CPW dimensions of conductor/gap are 6m/3m and resonant frequencies are designed to fall between 5.5-6.5 GHz. The narrow conductor and gap widths are chosen to increase the electric field participation at surfaces and interfaces to study their contribution to microwave loss Gao (2008); McRae et al. (2020). The electric field energy participation ratio of the metal/substrate interface for this CPW geometry is estimated to be between 0.4-0.5% based on the works of others Calusine et al. (2018); Kopas et al. (2022) where a 2nm thick interfacial layer was assumed. The electric field participation at the surfaces and interfaces of CPW resonators is more thoroughly discussed in other works McRae et al. (2020); Kopas et al. (2022); Calusine et al. (2018); Crowley et al. (2023). Notably, in order to compute such a participation ratio, assumptions must be made about the extent and dielectric properties of interfacial regions which may not be well understood.
Clean CPW resonator chips were wire-bonded with aluminum wire and packaged in a 2-port package constructed of gold-plated copper for microwave measurements in a dilution refrigerator with a base temperature of 35mK. The attenuation and amplification setup used is similar to that presented by McRae et al.McRae et al. (2020). Microwave loss of the resonators was characterized using a vector network analyzer (VNA) by fitting the power dependent spectra of the resonators to the diameter correction method (DCM)Khalil et al. (2012); McRae et al. (2020) model to extract internal () and external () resonator quality factors. A Josephson parametric amplifierAumentado (2020) installed at the 35mK stage was used for all low-power measurements (below 1000 photon occupation in the cavities).
Internal quality factor of resonators is commonly measured as a function of drive power and temperature in order to extract two-level system (TLS) loss and characterize materialsMcRae et al. (2020); Crowley et al. (2023).For the worst performing films in this work (epitaxial Ta grown directly on Al2O3(0001)), resonator is markedly poor and changes by less than 25% with power in contrast with the other samples presented where can vary by as much as two orders of magnitude over a similar input power range. For this reason, internal quality factor at low power only is used as a performance metric to compare films. This avoids consideration of these saturable sources of loss which are generally of interest to the community, but are beyond the scope of this work, while focusing on the low power metric that is most relevant to transmon performance. More discussion of quality factor power dependence is included in Appendix E.
@@ III.1 Overview of Experiment
Temperature dependent resistivity measurements of the Ta films grown at high temperature suggest high quality films. of epitaxial Ta films were measured to be near reported bulk values with RRR > 30 for 100nm thick films which is included in Table 2.
@@
The same microfabrication process is used for all samples presented here and is shown to yield high quality CPW resonators for all Nb films, but only for Ta films grown at lower temperature as shown in Figure 2. Ta films grown at higher temperature are epitaxial and high-quality structurally, but yield lossy resonators, seemingly at odds with structural and DC electrical characterization results.
In light of these observations, subsequent samples were fabricated to test the Ta/sapphire interface. The details of this experiment are provided in Section III.3, and the results suggest that the epitaxial Ta/sapphire interface is the dominant source of microwave loss at both high and low power in the samples having low . While our experiments identify the interface as the source of loss and provide techniques shown to eliminate it, the microscopic mechanism of the loss is left as an open question.
@@ III.2 Temperature dependent growth of Nb and Ta on
Figure 1 shows XRD measurement results taken from Nb and Ta films grown at varying substrate temperature. Each peak indicates a crystalline phase oriented perpendicular to the substrate surface, along the growth direction. For Nb, growth at room temperature results in a polycrystalline film having both (110) and (111) oriented grains along the growth direction. Non-specular measurements were also performed which show that the (111) and (110) grains have in-plane texture. XRD pole figures and more discussion are included in Appendix C. For Nb films grown at elevated temperature, only epitaxial Nb(111) is observed with no notable changes to structure as the growth temperature is increased in the temperature range studied. Throughout this work, epitaxial implies that the grown film has a single crystallographic orientation both out-of-plane and in-plane which is aligned to the underlying substrate. Non-specular measurements on these films confirm the expected epitaxial relationshipMayer et al. (1992) with .
XRD results from the Ta films (Fig. 1(b)) are somewhat more complex owing to the tendency of Ta thin films to crystallize in both the BCC (-Ta) and tetragonal (-Ta) structures. Ta grown at room temperature results in a film consisting mostly of highly textured -Ta(001) with a detectable amount of -Ta. In addition to out-of-plane texture, the -Ta(001) was found to be textured in-plane. Non-specular measurements (not included) show three rotational domains of 4-fold symmetric -Ta(001) aligned in-plane to the substrate, consistent with the 12-fold symmetric RHEED pattern (not included) observed immediately following growth. Ta deposition at 380 resulted in a biphasic film having both textured -Ta(001) and epitaxial -Ta(111) components, while growth at temperatures above 500 resulted in only epitaxial -Ta(111) having the same structural relationship to the substrate as the epitaxial Nb(111) described previously.
Sheet resistance at 300K, RRR, and were extracted from temperature dependent resistance measurements for the Nb and Ta films and is included in Tables 1 and 2 respectively. For Nb films, RRR increases with growth temperature as the film structure changes from textured Nb(110) to epi-Nb(111), but is somewhat reduced for the highest growth temperatures.
@@
For the textured -Ta(001) film grown at 20, resistivity increased with decreasing measurement temperature (RRR<1) and was below the base temperature of the He-4 cryostat used in the measurement. Temperature dependent transport of the epitaxial -Ta(111) samples was consistent with high quality films. A RRR value of 29 was measured for the -Ta(111) grown at 520 which increased to 47 for the film grown at 630. The of the -Ta(111) films range from 4.24K-4.35K, near reported values of pure, bulk Ta (4.49K)Inaba (1980). For the epitaxial Ta films, RRR is improved with increasing thickness. Appendix A includes data from films not included in the main text, with the best 200nm thick epitaxial Ta film reaching RRR=78.
AFM measurements were also performed on all of the films studied. While changes in morphology were observed with changing growth temperature, all films were found to have root-mean-square (RMS) roughness values below 2nm and no strong correlation to transport properties was observed. RMS roughness values for both Nb and Ta are included in Tables 1 and 2 respectively. Additional atomic force micrographs and discussion is included in Appendix D.
@@
Figure 2 shows of CPW resonators plotted against growth temperature for both Ta and Nb films. To obtain the data shown, spectra were measured for at least 3 resonators on each chip at varying input power corresponding to to below photons in the CPW cavities. values taken below 0.5 photon occupation (at least 3 per resonator) were then averaged together and plotted in Fig. 2. Thus, each data point shown includes multiple measurements of low-power from 3-8 different resonators measured from the same wafer, with the minimum, maximum, mean, and standard deviation indicated. These low-power data may be considered as figures of merit that are representative of the material and reflect the degree of MW loss at the ultra-low electric field intensity where transmons operate. The number of resonators measured on each chip, number of low power measurements on each resonator, and the range of fit parameters obtained from these measurements may be found in the Supplementary Information for all samples included in the main text.
@@
Measurement results from the Ta resonators are more interesting and somewhat unexpected. Firstly, it was found that resonators from the -Ta films grown at room temperature are high performing, with values exceeding those made from Nb resonators grown using the same conditions. Resonators containing -Ta were also shifted to lower frequency, presumably due to the kinetic inductance of the higher resistivity -Ta film Zmuidzinas (2012).
@@
While the markedly lower of the two epi-Ta films is apparent in Fig. 2, variations between the other samples presented is less clear. Notably, several films in the data presented have maximum values falling in the range 600k-800k, suggesting that the resonator performance may be limited by something common to them all, which includes the sapphire surface exposed to the dry etch when the Nb or Ta was etched away, the sapphire itself, and the backside of the substrate. A detailed analysis of the power dependent quality factor data with more measurements to provide the needed statistics may clarify optimum growth temperature for these films, though that is beyond the scope of this work.